OpenStudy (anonymous):

Find the distance between points (-3, 8) and (-5, 1). Find the distance between points (-4, -6) and (5, -1).

OpenStudy (anonymous):

The formula for calculating the distance between two points is \[\sqrt{(x_1-x_2)^2+(y_1-y_2)^2} \] So for the first two points you have \[\sqrt{(-3-(-5))^2+(8-1)^2}= \sqrt{(2)^2+(7)^2}= \sqrt{4+49}= \sqrt{53} \] You should be able to do the other one

OpenStudy (kymber):

The distance formula is  So plug in the points ;D